from the top of a building the angles of depression of the top of the base of a lamppost, located a distanse away were measured and found to be 58 degrees and 75 degrees,respectively if the building is 90 feet high.Find the height of the lamp post
1
Expert's answer
2012-07-26T10:52:05-0400
Distanse to lamp post is w=90/tan75
Also tan58=(90-h)/w, where h is height

So h=90(1-tan58/tan75)=-51.407